2.     Consider three functions f, g, and h (from R to R+). Let (x) = x2, g(x) = 2x and h(x) = log2(x)

a)  Evaluate f ◦ g(0)

b)  Evaluate h ◦ g(5)

c)  Evaluate h ◦ f(0)

d)  Give a mathematical expression for h ◦ g

e)  Give a mathematical expression for h ◦ g ◦ f

1
Expert's answer
2021-11-21T14:15:36-0500

a)

"f \u25e6 g(x)=(2^x)^2"

"f \u25e6 g(0)=(2^0)^2=1"


b)

"h \u25e6 g(x)=log_2(2^x)=x"

"h \u25e6 g(5)=5"


c)

"h \u25e6 f(x)=log_2(x^2)"

"h ◦ f(0)=log_2(0)=-\infin"


d)

"h \u25e6 g=log_2(2^x)=x"


e)

"h \u25e6 g \u25e6 f=log_2(2^{x^2})=x^2"